# Attributing the point symmetric structure of core-collapse supernova remnant N132D to the jittering jets explosion mechanism

### v1

@@ Abstract
Studies in the last two years have been discussing two competing core-collapse supernova (CCSN) explosion mechanisms, the delayed neutrino explosion mechanism (e.g., Andresen et al. 2024; Boccioli & Fragione 2024; Burrows et al. 2024; Janka & Kresse 2024; van Baal et al. 2024; Wang & Burrows 2024; Bamba et al. 2025; Boccioli et al. 2025; Eggenberger Andersen et al. 2025; Huang et al. 2025; Imasheva et al. 2025; Laplace et al. 2025; Maltsev et al. 2025; Maunder et al. 2024; Müller et al. 2025; Nakamura et al. 2025; Sykes & Müller 2025; Janka 2025; Paradiso & Coughlin 2025; Wang & Burrows 2025) and the jittering jets explosion mechanism (JJEM), which is the frame of the present study.
@@
The JJEM asserts that the primary explosion mechanism of the majority, and likely all, CCSNe involves pairs of jets (e.g., Soker 2010; Papish & Soker 2011). The magnetorotational explosion mechanism is much older, but significantly differs from the JJEM. In the magnetorotational explosion mechanism, one pair of jets along a fixed axis explodes the star. To maintain a fixed angular momentum axis along which the accretion disk launches a pair of jets, the progenitor of the CCSN must be rapidly rotating (e.g., Shibagaki et al. 2024; Zha et al. 2024; Shibata et al. 2025 and references to much older papers therein). The requirement for a rapidly rotating pre-collapse core makes this mechanism rare. Therefore, the magnetorotational explosion mechanism attributes most CCSNe to the neutrino-driven mechanism, and I classify it as part of the neutrino-driven mechanism.
@@
According to the JJEM, pairs of jittering jets that intermittent accretion disks (or belts) around the newly born neutron star (NS) launch on a timescales of explode all CCSNe (for a list of the parameters of the JJEM see Soker 2024a); time is measured from shock bounce, i.e., the formation time of the shock at the newly born NS. In rare cases when the pre-explosion core is rapidly rotating, the jittering of the pairs of jets is at very small angles around the pre-explosion angular momentum axis. Practically, it is a fixed-axis explosion, some of which end with a black hole remnant. The variations in the axes of the pairs of jittering jets might be fully or partially stochastic. The two jets in a pair might substantially differ in their opening angle and power (e.g., Bear et al. 2024), and might be at an angle smaller than , namely, not exactly opposite (e.g., Shishkin et al. 2025).
@@
Angular momentum fluctuations in the convection zones of the collapsing core that instabilities above the NS (for instabilities turbulence in the gain region, see, e.g., Abdikamalov et al. 2016; Kazeroni & Abdikamalov 2020; Buellet et al. 2023) amplify cause the stochastic variations in the angular momentum axis directions of the accreted gas that launches the jets along these axes (e.g., Gilkis & Soker 2014, 2016; Shishkin & Soker 2021, 2023; Wang et al. 2024). Neutrino heating plays a role by adding energy to the jets (Soker, 2022a). However, most of the energy originates from the jittering jets, which serve as the primary driver of the explosion.
In a recent review (Soker, 2024b), I summarized the state of debate between these two alternative explosion mechanisms as of 2024. The neutrino-driven mechanism encounters problems and difficulties that the community often overlooks. I list four issues.
@@
(2) Some modeling of CCSNe deduce explosion energies of (e.g., Moriya et al. 2025), more than the neutrino-driven mechanism can supply. Many magnetar models of superluminous CCSNe require explosion energies of that imply explosion by jets (e.g., Soker & Gilkis 2017; Kumar 2024), a conclusion overlooked by most magnetar modelers.
(3) The neutrino mechanism predicts that some massive stars fail to explode, in contradiction with studies that suggest that there is only a small or no population of failed CCSNe (e.g., Byrne & Fraser 2022; Strotjohann et al. 2024; Beasor et al. 2024; Healy et al. 2024). Boccioli et al. (2025) suggested that the neutrino-driven mechanism can overcome this difficulty.
(4) The most severe and challenging problem the neutrino-driven mechanism encounters is explaining the point-symmetric morphologies of CCSN remnants (CCSNRs; Section 2). This problem likely rules out the neutrino-driven mechanism as the primary mechanism of the explosion. On the other hand, the JJEM predicts that many CCSNRs exhibit point-symmetric morphologies, as recent three-dimensional hydrodynamical simulations demonstrate (Braudo et al., 2025). For this reason, identifying point-symmetric CCSNRs is of high significance in determining the explosion mechanism. In Section 3, I claim that SNR N132D is another CCSNR with a point-symmetric morphology that supports the JJEM.
@@ 2 Identifying point-symmetry by eyes
The pairs of opposite jets that explode the star in the JJEM might shape pairs of structural features in the descendant CCSNR. I emphasize that in most CCSNe these jets are not relativistic (Guetta et al. (2020) suggest that most CCSNe have no relativistic jets, contrary to gamma ray bursts that do have, e.g., Izzo et al. 2019; Abdikamalov & Beniamini 2025). Only a small number of pairs leave imprints on the descendant CCSNR; the other explode the core, lose their symmetry, and leave no clear geometrical imprint. Because the symmetry axes of the pairs of jets change their directions from one jet-launching episode to the next, the CCSNR’s pairs of structural features do not share the same axis, resulting in a point-symmetric morphology. The opposite structural features might include dense clumps, dense elongated structures termed filaments, bubbles, which are faint structures closed and encircled by a brighter rim, lobes, which are bubbles with partial rims, and ears, which are protrusions from the main CCSNR shell with decreasing cross-sections away from the center.
@@
Studies identified point-symmetric morphologies in 14 CCSNRs and discussed these in the frame of the JJEM. Some of these CCSNRs possess clear point-symmetrical morphological features, like N63A (Soker, 2024c), the Vela CCSNR (Soker 2023; Soker & Shishkin 2025), Cassiopeia A (Bear & Soker, 2025), and the Crab Nebula (Shishkin & Soker, 2024), and others with less secure point-symmetric morphologies, like CTB 1 (Bear & Soker, 2023). Other CCSNRs and the studies that attributed their morphologies to the JJEM are SNR 0540-69.3 (Soker, 2022b), the Cygnus Loop (Shishkin et al., 2024), SN 1987A (Soker, 2024d, e), G321.3–3.9 (Soker, 2024f; Shishkin & Soker, 2025), G107.7-5.1 (Soker, 2024f), W44 (Soker, 2024g), Puppis A (Bear et al., 2024), SNR G0.9+0.1 (Soker, 2025), and S147 Shishkin et al. 2025.
@@
Several processes are likely to disrupt the point-symmetric morphology, such as instabilities and the NS natal kick velocity that occur during the explosion, as well as certain post-explosion processes. The latter include interaction with a circumstellar material (CSM) lost by the CCSN progenitor (e.g., Chiotellis et al. 2021, 2024; Velázquez et al. 2023; Meyer et al. 2022, 2024b), interaction with the interstellar medium (ISM; e.g., Wu & Zhang 2019; Yan et al. 2020; Lu et al. 2021; Meyer et al. 2024a), pulsar wind nebula if it exists, and heating processes such as reverse shock and radioactive decay. In many CCSNRs, the processes that smear the point-symmetric structures induced by the explosion make the identification of point-symmetric morphologies challenging or even impossible.
@@
In this study of the SNR N132D and the earlier studies of point-symmetric CCSNRs, I have been utilizing my over thirty years of experience in classifying planetary nebulae by inspecting their jet-shaped morphologies. The visual-inspection classification of morphologies is a powerful common practice in classifying planetary nebulae (e.g., Balick 1987; Chu et al. 1987; Sahai et al. 2007, 2011), and AGN jets (e.g., Horton et al. 2025), and has led to significant breakthroughs, particularly in establishing the major role of binary interaction in the shaping of planetary nebulae, and the role that jets play in many of these interacting binaries.
@@
Although the visual-inspection method might seem completely subjective because it is not quantitative, it is not; it has a large objective component. Humans possess a remarkable ability to critically examine symmetry and identify departures from it by inspection alone, as facial symmetry is a key indicator of mate quality (e.g., Rhodes 2006; Pinheiro et al. 2023 and reference therein;). In particular, our ability to identify fluctuating asymmetries, which are non-directional (random) deviations from perfect symmetry in bilaterally paired traits, is crucial as fluctuating asymmetries tend to reflect health problems (e.g., Rhodes 2006). The majority of authors (and referees) made one of the most important decisions of their lives, if not the most important, i.e., choosing a mate, primarily through visual inspection, particularly by looking for symmetry and ruling out large fluctuating asymmetries. Recognition of symmetry is relatively automatic and consistent across cultures (e.g., Rhodes 2006).
@@ 3 The point symmetric structure of SNR N132D
Numerous papers study the structure and morphology of SNR N132D (e.g., Lasker 1978, 1980; Hughes 1987; Blair et al. 1994; Dickel & Milne 1995; Morse et al. 1995, 1996; Blair et al. 2000; Behar et al. 2001; Tappe et al. 2006, 2012; Xiao & Chen 2008; Bamba et al. 2018; Law et al. 2020; Sharda et al. 2020; Suzuki et al. 2020; Banovetz et al. 2023; Rho et al. 2023; XRISM Collaboration et al. 2024a; Foster et al. 2025; Okada et al. 2025). None of these studies examined the point-symmetric morphology and structure in the context of the JJEM, which is the goal of this study. Some studies have mentioned a possible jet as an explanation for the runaway knot in the southeast (e.g., Vogt & Dopita 2011), which I will refer to later.
Morse et al. (1995) first identified a point-symmetric structure of the fast O-rich ejecta of SNR N132D (Section 3.2); they did not term it a point-symmetric structure and did not connect it to jet shaping. Before turning to their results, I visually inspect some X-ray images to identify a global point-symmetric morphology (Section 3.1). In Section 3.3, I discuss the three-dimensional (3D) structure from Vogt & Dopita (2011) and Law et al. (2020) in the frame of the point symmetric morphology and the JJEM.
@@ 3.1 The X-ray point-symmetric morphology
The recent study by Foster et al. (2025) that maps iron emission, as I present in panel a of Figure 1, allows me to identify a symmetry axis, the short-symmetry axis of SNR N132D, extending from the northwest ear (NW ear) and through the center of the iron emission; this is the dashed-pale-blue line in panel a. The iron emission Fe K is concentrated to the southeast. XRISM Collaboration et al. (2024b) find that the Fe Ly emission from SNR N132D is redshifted with a bulk velocity of . In panel b of Figure 1, I present an X-ray image from the Chandra site. This X-ray image shows three dents (bent inwards) on the outer boundary of the SNR, which I have marked with three yellow arrows. A fourth dent is inside the boundary (yellow-dashed arrow). I connect two pairs of dents with two solid red lines. The solid-orange double-sided arrow is the short-symmetry axis, i.e., at the same location as the dashed line on panel a of Figure 1. I added a double-sided arrow (dashed orange) for orientation: it is perpendicular to the short-symmetry axis and at the same length; the two double-sided arrows intersect at their centers. As immediately seen, SNR N132D is elongated in the northeast. The southwest end of the perpendicular double-sided arrow does not reach the boundary of the SNR, implying that the SNR is also elongated to the southwest.
@@
Banovetz et al. (2023) measured the proper motion of the oxygen-rich ejecta of SNR N132D from two HST observations 16 years apart. They determined the center of expansion (CoE); in panel c of Figure 1, which I adapted from Banovetz et al. (2023), they mark this center with a yellow plus symbol. They also mark the two centers that Morse et al. (1995) identified, which I refer to in Section 3.2. I added to panel c the intersections of the two double-sided arrows and of the two red lines on panel b of Figure 1. The Chandra X-ray image of SNR N132D in panel d of Figure 1, adapted from Borkowski et al. (2007), shows the global structure of SNR N132D, including the centers of different structures from panel c, the structures from panel b, and the contours of the oxygen-rich material near the center (more on the oxygen-rich ejecta in Section 3.2 and 3.3).
@@
While I based the short-symmetry axis, which is the double-sided arrows from southeast to the NW ear, on structural features, i.e., the NW ear and the iron-rich zone from Foster et al. (2025), the dashed-orange double-sided arrow is simply a perpendicular line with the same length and through the center of the short-symmetry axis. A key feature for identifying point-symmetric morphology is that this perpendicular line aligns with the longest structure of the oxygen-rich ejecta near the center, indicating a clear association. However, I can also draw a line from a small ear on the southwest and through the center of the short-symmetry axis; this is the dotted-orange double-sided arrow. The southwest end of the dotted-orange double-sided arrow touches the edge of that ear, and its center is at the center of the short-symmetry axis. The dashed and dotted orange double-sided arrows are to each other.
@@
I attribute the long extension of SNR N132D to a pair of opposite jets (or more). The angle of reflects the uncertainty in the location of the jets’ axis. Alternatively, the two jets were not exactly opposite to each other, but rather had an angle of between them. Such a ‘bent symmetry’, i.e., the two opposite sides are close to being at but not exactly, is observed in many planetary nebulae. Shishkin et al. (2025) suggest such bent opposite jets for CCSNR S147. I mark with solid arrows on panel d of Figure 1 the two jets I propose for this possibility.
@@
The main result of this section is the identification of two axes that describe the large-scale structure, the short-symmetry axis and the perpendicular axis that extend along the lone dimension of SNR N132D. I turn to show, Sections 3.2 and 3.3, that these axes correlate with those of the inner structure of the oxygen-rich ejecta.

### v2

@@ Abstract
Studies in the last two years have been discussing two competing core-collapse supernova (CCSN) explosion mechanisms, the delayed neutrino explosion mechanism (e.g., Andresenetal2024; BoccioliFragione2024; Burrowsetal2024kick; JankaKresse2024; vanBaaletal2024; WangBurrows2024; Bambaetal2025CasA; Bocciolietal2025; EggenbergerAndersenetal2025; Huangetal2025; Imashevaetal2025; Laplaceetal2025; Maltsevetal2025; Maunderetal2025; Morietal2025; Mulleretal2025; Nakamuraetal2025; SykesMuller2025; Janka2025; ParadisoCoughlin2025; Vinketal2025; WangBurrows2025) and the jittering jets explosion mechanism (JJEM), which is the frame of the present study.
@@
The JJEM asserts that the primary explosion mechanism of the majority, and likely all, CCSNe involves pairs of jets (e.g., Soker2010; PapishSoker2011). The magnetorotational explosion mechanism is much older, but significantly differs from the JJEM. In the magnetorotational explosion mechanism, one pair of jets along a fixed axis explodes the star. To maintain a fixed angular momentum axis along which the accretion disk launches a pair of jets, the progenitor of the CCSN must be rapidly rotating (e.g., Shibagakietal2024; ZhaMullerPowell2024; Shibataetal2025 and references to much older papers therein). The requirement for a rapidly rotating pre-collapse core makes this mechanism rare. Therefore, the magnetorotational explosion mechanism attributes most CCSNe to the neutrino-driven mechanism, and I classify it as part of the neutrino-driven mechanism.
@@
According to the JJEM, pairs of jittering jets that intermittent accretion disks (or belts) around the newly born neutron star (NS) launch on a timescales of explode all CCSNe (for a list of the parameters of the JJEM see Soker2025Learning); time is measured from shock bounce, i.e., the formation time of the shock at the newly born NS. In rare cases when the pre-explosion core is rapidly rotating, the jittering of the pairs of jets is at very small angles around the pre-explosion angular momentum axis. Practically, it is a fixed-axis explosion, some of which end with a black hole remnant. The variations in the axes of the pairs of jittering jets might be fully or partially stochastic. The two jets in a pair might substantially differ in their opening angle and power (e.g., Bearetal2025Puppis), and might be at an angle smaller than , namely, not exactly opposite (e.g., Shishkinetal2025S147).
@@
Angular momentum fluctuations in the convection zones of the collapsing core that instabilities above the NS (for instabilities turbulence in the gain region, see, e.g., Abdikamalovetal2016; KazeroniAbdikamalov2020; Buelletetal2023) amplify cause the stochastic variations in the angular momentum axis directions of the accreted gas that launches the jets along these axes (e.g., GilkisSoker2014; GilkisSoker2016; ShishkinSoker2021; ShishkinSoker2023; WangShishkinSoker2024). Neutrino heating plays a role by adding energy to the jets (Soker2022nu). However, most of the energy originates from the jittering jets, which serve as the primary driver of the explosion.
In a recent review (Soker2024UnivReview), I summarized the state of debate between these two alternative explosion mechanisms as of 2024. The main challenge of the JJEM is to demonstrate with three-dimensional numerical simulations the formation of intermittent accretion disks during the explosion process. Such simulations should be with very high resolution, include magnetic fields, and the correct pre-collapse core convection. The next step of launching the jets by these intermittent accretion disks seems to be beyond the present capabilities of numerical simulations Soker2025Learning. The neutrino-driven mechanism encounters problems and difficulties that the community often overlooks. I list four issues.
@@
(2) Some modelings of CCSNe deduce explosion energies of (e.g., Moriyaetal2025), more than the neutrino-driven mechanism can supply. Many magnetar models of superluminous CCSNe require explosion energies of that imply explosion by jets (e.g., SokerGilkis2017; Kumar2025), a conclusion overlooked by most magnetar modelers.
(3) The neutrino mechanism predicts that some massive stars fail to explode, in contradiction with studies that suggest that there is only a small or no population of failed CCSNe (e.g., ByrneFraser2022; StrotjohannOfekGalYam2024; BeasoretalLuminosity2025; Healyetal2025). Bocciolietal2025 suggested that the neutrino-driven mechanism can overcome this difficulty.
(4) The most severe and challenging problem the neutrino-driven mechanism encounters is explaining the point-symmetric morphologies of CCSN remnants (CCSNRs; Section 2). This problem likely rules out the neutrino-driven mechanism as the primary mechanism of the explosion. On the other hand, the JJEM predicts that many CCSNRs exhibit point-symmetric morphologies, as recent three-dimensional hydrodynamical simulations demonstrate (Braudoetal2025). For this reason, identifying point-symmetric CCSNRs is of high significance in determining the explosion mechanism. In Section 3, I claim that SNR N132D is another CCSNR with a point-symmetric morphology that supports the JJEM.
@@ 2 Identifying point-symmetry by eyes
The pairs of opposite jets that explode the star in the JJEM might shape pairs of structural features in the descendant CCSNR. I emphasize that in most CCSNe these jets are not relativistic (Guettaetal2020 suggest that most CCSNe have no relativistic jets, contrary to gamma ray bursts that do have, e.g., Izzoetal2019; AbdikamalovBeniamini2025). Only a small number of pairs leave imprints on the descendant CCSNR; the other explode the core, lose their symmetry, and leave no clear geometrical imprint. Because the symmetry axes of the pairs of jets change their directions from one jet-launching episode to the next, the CCSNR’s pairs of structural features do not share the same axis, resulting in a point-symmetric morphology. The opposite structural features might include dense clumps, dense elongated structures termed filaments, bubbles, which are faint structures closed and encircled by a brighter rim, lobes, which are bubbles with partial rims, and ears, which are protrusions from the main CCSNR shell with decreasing cross-sections away from the center.
@@
Studies identified point-symmetric morphologies in 14 CCSNRs and discussed these in the frame of the JJEM. Some of these CCSNRs possess clear point-symmetrical morphological features, like N63A (Soker2024CounterJet), the Vela CCSNR (Soker2023SNRclass; SokerShishkin2025Vela), Cassiopeia A (BearSoker2025), and the Crab Nebula (ShishkinSoker2025Crab), and others with less secure point-symmetric morphologies, like CTB 1 (BearSoker2023RNAAS). Other CCSNRs and the studies that attributed their morphologies to the JJEM are SNR 0540-69.3 (Soker2022SNR0540), the Cygnus Loop (ShishkinKayeSoker2024), SN 1987A (Soker2024NA1987A; Soker2024Keyhole), G321.3–3.9 (Soker2024CF; ShishkinSoker2025G321), G107.7-5.1 (Soker2024CF), W44 (Soker2024W44), Puppis A (Bearetal2025Puppis), SNR G0.9+0.1 (Soker2025G0901), and S147 Shishkinetal2025S147. In some CCSNRs, the dust also presents point-symmetric morphologies (Soker2025Dust).
@@
Several processes are likely to disrupt the point-symmetric morphology, such as instabilities and the NS natal kick velocity that occur during the explosion, as well as certain post-explosion processes. The latter include interaction with a circumstellar material (CSM) lost by the CCSN progenitor (e.g., Chiotellisetal2021; ChiotellisZapartasMeyer2024; Velazquezetal2023; Meyeretal2022; MeyerDetal2024), interaction with the interstellar medium (ISM; e.g., Wuetal2019; YanLuetal2020; LuYanetal2021; MeyerMelianietal2024), pulsar wind nebula if it exists, and heating processes such as reverse shock and radioactive decay. In many CCSNRs, the processes that smear the point-symmetric structures induced by the explosion make the identification of point-symmetric morphologies challenging or even impossible.
@@
In this study of the SNR N132D and the earlier studies of point-symmetric CCSNRs, I have been utilizing my over thirty years of experience in classifying planetary nebulae by inspecting their jet-shaped morphologies. The visual-inspection classification of morphologies is a powerful common practice in classifying planetary nebulae (e.g., Balick1987; Chuetal1987; Sahaietal2007; Sahaietal2011), and AGN jets (e.g., Hortonetal2025), and has led to significant breakthroughs, particularly in establishing the major role of binary interaction in the shaping of planetary nebulae, and the role that jets play in many of these interacting binaries.
@@
Although the visual-inspection method might seem completely subjective because it is not quantitative, it is not; it has a large objective component. Humans possess a remarkable ability to critically examine symmetry and identify departures from it by inspection alone, as facial symmetry is a key indicator of mate quality (e.g., Rhodes2006; Pinheiroetal2023 and reference therein;). In particular, our ability to identify fluctuating asymmetries, which are non-directional (random) deviations from perfect symmetry in bilaterally paired traits, is crucial as fluctuating asymmetries tend to reflect health problems (e.g., Rhodes2006). The majority of authors (and referees) made one of the most important decisions of their lives, if not the most important, i.e., choosing a mate, primarily through visual inspection, particularly by looking for symmetry and ruling out large fluctuating asymmetries. Recognition of symmetry is relatively automatic and consistent across cultures (e.g., Rhodes2006).
@@ 3 The point symmetric structure of SNR N132D
Numerous papers study the structure and morphology of SNR N132D (e.g., Lasker1978; Lasker1980; Hughes1987; Blairetal1994; DickelMilne1995AJ; Morseetal1995; Morseetal1996; Blairetal2000; Beharetal2001; Tappeetal2006; Tappeetal2012; XiaoAndChen2008; Bambaetal2018; Lawetal2020; Shardaetal2020; Suzukietal2020; Banovetzetal2023; Rhoetal2023; Xrism2024N132D; Fosteretal2025; LongX2025; Okadaetal2025). None of these studies examined the point-symmetric morphology and structure in the context of the JJEM, which is the goal of this study. Some studies have mentioned a possible jet as an explanation for the runaway knot in the southeast (e.g., VogtDopita2011), which I will refer to later.
Morseetal1995 first identified a point-symmetric structure of the fast O-rich ejecta of SNR N132D (Section 3.2); they did not term it a point-symmetric structure and did not connect it to jet shaping. Before turning to their results, I visually inspect some X-ray images to identify a global point-symmetric morphology (Section 3.1). In Section 3.3, I discuss the three-dimensional (3D) structure from VogtDopita2011 and Lawetal2020 in the frame of the point symmetric morphology and the JJEM.
@@ 3.1 The X-ray point-symmetric morphology
The recent study by Fosteretal2025 that maps iron emission, as I present in panel (a) of Figure 1, allows me to identify a symmetry axis, the short-symmetry axis of SNR N132D, extending from the northwest ear (NW ear) and through the center of the iron emission; this is the dashed-pale-blue line in panel (a). The X-ray image from the Chandra site, panel (b) of Figure 1, reveals a clear ear in the northwest. The protrusion on the opposite side is small and not distinct enough to be identified as a separate structural feature, although the two dents are. The new image by Fosteretal2025 presents a clear opposite structure to the northwest ear: a zone extending from the center to the southeast: the yellow-colored zones in panel (a) of Figure 1 that depict strong Fe K emission. The line along the long axis of the strong Fe K emission in the southeast crosses to the other side into the northwest ear. This is the short symmetry axis. The bright zones of other emission lines might be in different regions than the bright Fe K region, e.g., the S He, Fe He, and Fe Ly in the XRISM maps (Xrism2024N132D); these maps have lower spatial resolution than the XMM-Newton maps of Fosteretal2025. I use the bright zone of the Fe K line from Fosteretal2025 because this zone occupies a well-defined region that is bound by the two dents that I mark on panel (b) of Figure 1.
@@
The iron emission Fe K is concentrated to the southeast. Xrism2024N132D find that the Fe Ly emission from SNR N132D is redshifted with a bulk velocity of . In panel (b) of Figure 1, I present an X-ray image from the Chandra site. This X-ray image shows three dents (bent inwards) on the outer boundary of the SNR, which I have marked with three yellow arrows. A fourth dent is inside the boundary (yellow-dashed arrow). I connect two pairs of dents with two solid red lines. The solid-orange double-sided arrow is the short-symmetry axis, i.e., at the same location as the dashed line on panel (a) of Figure 1. I added a double-sided arrow (dashed orange) for orientation: it is perpendicular to the short-symmetry axis and at the same length; the two double-sided arrows intersect at their centers. As immediately seen, SNR N132D is elongated in the northeast. The southwest end of the perpendicular double-sided arrow does not reach the boundary of the SNR, implying that the SNR is also elongated to the southwest.
@@
Banovetzetal2023 measured the proper motion of the oxygen-rich ejecta of SNR N132D from two HST observations 16 years apart. They determined the center of expansion (CoE); in panel (c) of Figure 1, which I adapted from Banovetzetal2023, they mark this center with a yellow plus symbol. They also mark the two centers that Morseetal1995 identified, which I refer to in Section 3.2. I added to panel (c) the intersections of the two double-sided arrows and of the two red lines on panel (b) of Figure 1. The Chandra X-ray image of SNR N132D in panel (d) of Figure 1, adapted from Borkowskietal2007, shows the global structure of SNR N132D, including the centers of different structures from panel (c), the structures from panel (b), and the contours of the oxygen-rich material near the center (more on the oxygen-rich ejecta in Section 3.2 and 3.3).
@@
While I based the short-symmetry axis, which is the double-sided arrows from southeast to the NW ear, on structural features, i.e., the NW ear and the iron-rich zone from Fosteretal2025, the dashed-orange double-sided arrow is simply a perpendicular line with the same length and through the center of the short-symmetry axis. A key feature for identifying point-symmetric morphology is that this perpendicular line aligns with the longest structure of the oxygen-rich ejecta near the center, indicating a clear association. However, I can also draw a line from a small ear on the southwest and through the center of the short-symmetry axis; this is the dotted-orange double-sided arrow. The southwest end of the dotted-orange double-sided arrow touches the edge of that ear, and its center is at the center of the short-symmetry axis. The dashed and dotted orange double-sided arrows are to each other.
@@
I attribute the long extension of SNR N132D to a pair of opposite jets (or more). The angle of reflects the uncertainty in the location of the jets’ axis. Alternatively, the two jets were not exactly opposite to each other, but rather had an angle of between them. Such a ‘bent symmetry’, i.e., the two opposite sides are close to being at but not exactly, is observed in many planetary nebulae. Shishkinetal2025S147 suggest such bent opposite jets for CCSNR S147. I mark with solid arrows on panel (d) of Figure 1 the two jets I propose for this possibility.
@@
The main result of this section is the identification of two axes that describe the large-scale structure, the short-symmetry axis and the perpendicular axis that extend along the long dimension of SNR N132D. I turn to show, Sections 3.2 and 3.3, that these axes correlate with those of the inner structure of the oxygen-rich ejecta.